OpenStudy (anonymous):

HELP!!! I"LL MEDAL!!!! x+2-3=3^2

OpenStudy (aaronandyson):

x + 2 - 3 = 3^2 x + 2 - 3 = 9 x - 1 = 9 x = 9 + 1 x = 10 You're correct ! :)

OpenStudy (anonymous):

thx